Please answer Q1 (A,B,C) Will Rate Thank You!

A) What is the equivalent resistance of 100 resistors each of 0.01Ω connected in parallel to an external circuit?

B) Calculate the decay constant (lambda) of radioactive materials of half-life 10 ns?

C) What is the electric field (E) produced by the source charge (+10 Coul.) at a distance of 0.01 m? (ke = 8.99 X 10^9 N·m^2/C^2)
Answer #1

A) The Equivalent Resistance of n resistors connected in parallel is given by

R_{eq}= \frac{R}{n}

Given:

R =0.01 \Omega

n = 100

R_{eq}= \frac{0.01 \Omega}{100}

Reg 1 x 10 -4\Omega

B) The Decay Constant of a Radioactive material is given by

\lambda = \frac{0.6931}{T_{\frac{1}{2}}}

Given

T1/2 = 10 x 10-9 s

\lambda = \frac{0.6931}{10*10^{-9}}

\lambda = 6.931 * 10^{7} s^{-1}

C) The Electric Field due to a point charge at a distance r is given by

  E = \frac{k q}{r^2}

Given

k = 8.99 x 109Nm2C-2

r = 0.01 m

q = 10 C

E = \frac{8.99 * 10^9 *10}{(0.01)^2}

  E = 8.99 * 10^1^4 NC^{-1}

Hope this Helps!!